OpenStudy (anonymous):

simplify the following rational expression: x 2 -4x-5 x-5

OpenStudy (anonymous):

can you clarify the expression, it is unclear

OpenStudy (anonymous):

is it \[\frac{x^2-4x-5}{x-5}\]?

OpenStudy (anonymous):

factor and cancel \[\frac{(x-5)(x+1)}{x-5}=\frac{\cancel{(x-5)}(x+1)}{\cancel{x-5}}=x+1\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

thanks god! someone finally figured it out!

OpenStudy (anonymous):

method is clear right? factor if you can, then cancel the common factors

OpenStudy (anonymous):

yep :) can u help me with another one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@katlin95 i helped u out, and i told u if what i was telling u was right relating to the equation was right and u would say, yes... but it was not :/

OpenStudy (anonymous):

@katlin95 u sould but ur question a bit more clear

OpenStudy (anonymous):

which expression represents 12 x 3 -6 x 2 +2x 2x in simplest form

OpenStudy (anonymous):

are those times signs, or \(x\)'s

OpenStudy (anonymous):

like the one above ^ its not clear. this usually happens when someone just copies and paste the question from online.

OpenStudy (anonymous):

A. 6 x 2 -3x B. 10 x 2 -4x C. 6 x 2 -3x+1 D. 10 x 2 -4x+1 they are x's:)

OpenStudy (anonymous):

@satellite73 ? where did u go?

OpenStudy (anonymous):

is it like this \[12 x^ 3 -6 x ^2 +2x\times 2x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

she kept telling me yes that i was correct also but at the end the equation i was talking about was nothing like how it was.

OpenStudy (anonymous):

ok then \[2x\times2x=4x^2\] so you have \[12x^3-6x^2+4x^2=12x^2-2x^2\]

OpenStudy (anonymous):

but that is not one of your choices, so i guess i read it wrong

OpenStudy (anonymous):

maybe you could rewrite it, use ^ for exponents, like x^2 for x squared

OpenStudy (anonymous):

@satellite73, no she is the one telling us the equation and writing it unclear

OpenStudy (anonymous):

or else post the actual question as an attachment